y′′′−2y′′−y′+2y=e4x


variation of paramaters


1
Expert's answer
2020-11-22T18:46:00-0500

Let us solve the differential equation "y'''\u22122y''\u2212y'+2y=e^{4x}" using the method of variation of paramaters. For this, firstly solve the characteristic equation "k^3-2k^2-k+2=0" of "y'''\u22122y''\u2212y'+2y=0". This characteristic eqution is equivalent to "k^2(k-2)-(k-2)=0" which is equivalent to "(k^2-1)(k-2)=0". Therefore, "k_1=1,k_2=-1,k_3=2." According to the method of variation of paramaters, the solution of "y'''\u22122y''\u2212y'+2y=e^{4x}" is

"y(x)=C_1(x)e^x+C_2(x)e^{-x}+C_3(x)e^{2x}."


Then we have the following system:


"\begin{cases}
C_1'(x)e^x+C_2'(x)e^{-x}+C_3'(x)e^{2x}=0\\
C_1'(x)e^x-C_2'(x)e^{-x}+2C_3'(x)e^{2x}=0\\
C_1'(x)e^x+C_2'(x)e^{-x}+4C_3'(x)e^{2x}=e^{4x}
\end{cases}"


which is equivalent to


"\begin{cases}
C_1'(x)e^x+C_2'(x)e^{-x}+C_3'(x)e^{2x}=0\\
2C_1'(x)e^x+3C_3'(x)e^{2x}=0\\
3C_3'(x)e^{2x}=e^{4x}
\end{cases}"


"\begin{cases}
C_2'(x)=-C_1'(x)e^{2x}-C_3'(x)e^{3x}\\
C_1'(x)=-\frac{3}{2}C_3'(x)e^{x}\\
C_3'(x)=\frac{1}{3}e^{2x}
\end{cases}"


"\begin{cases}
C_2'(x)=\frac{1}{2}e^{5x}-\frac{1}{3}e^{5x}\\
C_1'(x)=-\frac{1}{2}e^{3x}\\
C_3'(x)=\frac{1}{3}e^{2x}
\end{cases}"


"\begin{cases}
C_2'(x)=\frac{1}{6}e^{5x}\\
C_1'(x)=-\frac{1}{2}e^{3x}\\
C_3'(x)=\frac{1}{3}e^{2x}
\end{cases}"


"\begin{cases}
C_2(x)=\frac{1}{30}e^{5x}+C_2\\
C_1(x)=-\frac{1}{6}e^{3x}+C_1\\
C_3(x)=\frac{1}{6}e^{2x}+C_3
\end{cases}"


Therefore, the solution is the following:


"y(x)=(-\frac{1}{6}e^{3x}+C_1)e^x+(\frac{1}{30}e^{5x}+C_2)e^{-x}+(\frac{1}{6}e^{2x}+C_3)e^{2x}="

"=C_1e^x+C_2e^{-x}+C_3e^{2x}+\frac{1}{30}e^{4x}."